OpenStudy (anonymous):

suppose that a population of skunks lives on an isolated island is 500 skunks are present today and the birth rate is 65 per week and the death is 47 per week how many skunks to you predict will be on the island 105 days from now and what would a graph population versus time look for this population? Answer choices: 230 skunks; graphs shows a downward trend 550 skunks; graph shows a stable trend 770 skunks; graph shows an upward trend 432 skunks; graph shows a downward trend

OpenStudy (anonymous):

@Ashleyisakitty @hartnn @wolfe8

OpenStudy (anonymous):

Is the answer: 770 skunks; graph shows an upward trend?

OpenStudy (anonymous):

or would it be 550 skunks; graph shows a stable trend?

OpenStudy (wolfe8):

I would say since it's multiple choice, yes that once(without calculation). This is because the rate of birth is higher than the rate of death, meaning they will increase in number i.e there will be more coming to life than dying.

OpenStudy (wolfe8):

To calculate, first find how many number or weeks 105 days is. Then you know every week you will have 65+(-47) additional skunks. So you will add 500 to the multiple of number of weeks with addition each week. Does that make sense?

OpenStudy (anonymous):

Yes, so: \[65 + (-47) = 18\] \[105 \div 7 = 15 weeks\] What would I do next?

OpenStudy (wolfe8):

Then you find out how many skunks will be added in 15 weeks if each week you get 18

OpenStudy (anonymous):

So?: \[15 \times 18 = 270 + 500 = 770?\]

OpenStudy (wolfe8):

Yeppers :) And thanks for the nice testimony

OpenStudy (anonymous):

Thank you! I'm glad to give you a testimony :-)

OpenStudy (wolfe8):

Haha. Hope this helps you with future questions as well. Have a nice day.

OpenStudy (anonymous):

You, too!